Circular No. 7571

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V445 PUPPIS R. M. Wagner, LBT Observatory; G. Schwarz, Steward Observatory,

and S. G. Starrfield, Arizona State University, report: “Spectra (range 320– 960 nm, resolution 0.85 nm) of V445 Pup were obtained on Jan. 19.31 UT with the Bok 2.3-m telescope (+ Boller & Chivens spectrograph) on Kitt Peak. The spectrum has changed little since our previous report (IAUC 7556). In addition to the features present in our earlier spectrum, emission lines of Fe II 325.5- and 327.7-nm; O I 747.7-, 777.4-, 822.7-, 844.6-, 882.0-, and 926.4-nm; Mg I 383.5-, 517.8-, 552.8-, 880.7-, and 892.4-nm; Mg II 789.0-, 823.2-, and 922.6-nm; N I 821.2-, 869.2-, 905.0-, and 939.6-nm; C I 833.5-, 908.8-, and 940.6-nm; and Ca II 393.3-, 396.8-, 849.8-, 854.2-, and 866.2-nm are also present, superposed on a blue continuum. Some of these features exhibit P-Cyg profiles. Hydrogen lines arising from both the Balmer and Paschen series, as well as He II 468.6-nm, are very weak or absent (cf. IAUC 7556, 7559). A weak absorption line at 578.0 nm with an equivalent width of 0.12 nm may be partially interstellar, implying E(B−V) ≤ 0.8 mag. Apart from the absence of the Balmer lines, the overall spectrum resembles ‘Fe II’-class novae early in their evolution (Williams et al. 1994, Ap.J. Suppl. 90, 297). However, the apparent presence of He I (IAUC 7556, 7559) and a blue continuum, appearing simultaneously with permitted lines arising from neutral and low-ionization species, is not typical of these novae at such an early evolutionary phase.”

SUPERNOVA 2001I IN UGC 2836 A. V. Filippenko and R. Chornock, University of California, Berkeley,

report that inspection of CCD spectra (range 330–1000 nm) of SN 2001I (cf. IAUC 7563), obtained on Jan. 23 UT with the Shane 3-m reflector at Lick Observatory, reveals that the object is a type-IIn supernova. Strong hydrogen Balmer (and much weaker He I) emission lines are superposed on a nearly featureless, blue continuum.

SUPERNOVA 2001G IN MCG +08-17-43 Corrigendum. On IAUC 7569, for SUPERNOVA 2001G IN IC 677

read SUPERNOVA 2001G IN MCG +08-17-43

2001 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green